OpenStudy (anonymous):

Find the integral of (sin(pi*z^2)+cos(pi*z^2))/(z-1)(z-2) with respect to absolute z=3. and Find the integral of (e^2z)/(z+1)^4 with respect to absolute z=3.

OpenStudy (experimentx):

you mean radius 3 ??

OpenStudy (experimentx):

|z| = 3?

OpenStudy (anonymous):

Yes.

OpenStudy (experimentx):

make use of residue theorem ... this case both z=1,2 lies both inside the radii |z| = 3

OpenStudy (experimentx):

the first one is easy .. .i guess.

OpenStudy (experimentx):

for second one ... \[ \lim_{z \rightarrow -1} {1 \over 4!} {d\over dz}(z+1)^4{e^{2z}\over (z+1)^4}\]

OpenStudy (experimentx):

Oops!! \[ \lim_{z \rightarrow -1} {1 \over 4!} {d^3\over dz^3}(z+1)^4{e^{2z}\over (z+1)^4}\]

OpenStudy (experimentx):

that would be the residue for second. The itegral would be 2 \pi i Residue( f(z) ) which you can calculate as above.

OpenStudy (anonymous):

WIth the problem of sin(z) =2, the steps I followed was to express (sin (x)= (e^z-e^(-1))/2i = 2 and (e^z-e^(-1) = 4i ending up with e^x * sin y =2. Is that the correct way to go or is there an alternative.

OpenStudy (experimentx):

(e^z-e^(-z) = 4i (e^z)^2 - 4i e^z- 1 = 0 x^2 - 4i x - 1 = 0 x = 4i +- sqrt((4i)^2 + 4))/2 e^z = .... z = ln(....)

OpenStudy (anonymous):

Many thanks. In the proof that all values of (1-i)^(isqrt{2}) lie on a straight line can you assist with the steps to follow?

OpenStudy (experimentx):

if it has more than one root in [0, 2pi] then it should be false. If it has only one root in .. then [0, 2pi], then it would be branches of it ... which would mean that they all lie in straight line.

OpenStudy (experimentx):

I am not so sure .... this is just mu hunch ... you better try asking people on http://math.stackexchange.com